rxcklesskaisher:

A farmer has decided to divide his land area in half in order to plant soy and corn. Calculate the area of the entire area so he knows how much soil is needed. Each bag of soil covers 15 square yards. How many bags should the farmer purchase? A) 1 bag B) 4 bags C) 3 bags D) 2 bags

rxcklesskaisher:

1 attachment
AZ:

So first let's calculate the area of the plot we have a parallelogram the base is 2.5 + 6.5 what is the total base? and the height is 6 remember area = base * height

rxcklesskaisher:

i was afk

rxcklesskaisher:

@AZ

AZ:

yeah so what's the base? what is 2.5 + 6.5 = ?

rxcklesskaisher:

9

AZ:

okay so our base is 9 and our height from the image was 6 area = base * height area = 9 * 6 what is the area ?

rxcklesskaisher:

54

AZ:

Good so the area is 54 square yards and we know that 1 bag of soil covers 15 square yards so how many bags do we need to cover 54 square yards? just divide 54 / 15 =

rxcklesskaisher:

3.6

AZ:

so how many bags would he need? remember you have to round up

rxcklesskaisher:

4

AZ:

and that's your answer!